OpenStudy (anonymous):

What is the solution to the equation log(2)x - log(2)4 = 5 ?

Parth (parthkohli):

\(\log_24 = 2\). Agree?

OpenStudy (anonymous):

yes, thank you!

Parth (parthkohli):

Nope, not done yet. \( \color{Black}{\Rightarrow \log_2x - \log_24 = 5 \Longrightarrow \log_2x - 2 = 5}\) \( \color{Black}{\Rightarrow \log_2x = 7}\)

Parth (parthkohli):

Can you convert \(\log_2x = 7\) to an exponential form?

OpenStudy (anonymous):

Another way to skin the cat is to recognize: \[\log_{2} x - \log_{2} y = \log_{2}(x / y)\] So you have \[\log_{2}(x/4) = 5 \rightarrow 2^5 = x/4 \rightarrow 32 = x/4 -> ?\]